An amount of charge is distributed uniformly throughout a spherical shell that has a diameter 0.52m. At the center, there is a point charge q = 70 μC. Find the electric field outside the shell.

1
Expert's answer
2022-04-13T09:20:52-0400

"E=\frac{4kq}{d^2}=5.3~\frac{MN}C."